OpenStudy (anonymous):

Write an equation for the translation of y=4/x that has the asymptotes x=7 and y =6

OpenStudy (amistre64):

thats another wayto say: center it at (7,6)

OpenStudy (amistre64):

y=4/x is centered at 0,0 we want the center to be 7,6; lets move this by -7,-6 to line up the center with 0,0 but in order to move one point, we have to move all the (x,y) points by the same amount: (x-7, y-6) y=4/x -> y-6=4/(x-7)